# Solving Empirical Bayes via Transformers

### v1

@@ 7 Conclusion and Future Work
We have demonstrated the ability of transformers to learn EB-Poisson via in-context learning. This was done by evaluating pre-trained transformers on synthetic data of unseen distribution and sequence length, and compared against baselines like the NPMLE. In this process, we showed that transformers can achieve decreasing regret as the sequence length increases. On the real datasets, we showed that these pre-trained transformers can outperform classical baselines in most cases.
One future direction will be to extend our work to multi-dimensional input, as discussed in [JPTW23] (Section 1.3), [JPW22] (Section 6). We believe that the transformers would still be able to learn the ‘context’ of the inputs in multi-dimensional settings. On the other hand, the $g$-modelling methods like the NPMLE can take $n^{\Theta(d)}$ inference time, which makes it not scalable. In addition, given that the focus of this work is on Poisson-EB, one natural direction is to extend it to the normal-means model [JZ09]. On the theoretical front, the expressibility and limitations of the transformers can be further studied, including settings where the model dimension is bounded. Finally, given that the focus has been studying transformers trained and evaluated on priors with compact support ($[0,50]$ in our case), we plan to study further the behavior of transformers on priors with unbounded support (akin to how we did in one of the studies in Section 5.1).
@@ Acknowledgements
This work was supported in part by the MIT-IBM Watson AI Lab and the National Science Foundation under Grant No CCF-2131115. Anzo Teh was supported by a fellowship from the Eric and Wendy Schmidt Center at the Broad Institute.
@@ A.1 Worst-case prior and Gold-Standard Estimator
Let $A$ be a compact subset of $\mathcal{R}$. Then the worst-case prior $\pi_{!,A}$ is defined as
@@
One motivation for using the worst case prior is that the Bayes estimator is considered the “gold standard” which minimizes the maximum-possible MSE across all priors supported on $A$. A concrete statement can be found in the following lemma.
@@
Let $\hat{\theta}_{\pi}$ be the Bayes estimator to a prior $\pi$. and let $A$ be any compact subset of the reals. Then the least favorable prior $\pi_{!,A}$ of $A$ satisfies the following:
@@
For any compact subset $A$ of the reals, we have
@@
achieved by the Bayes estimator $f_{\pi_{!,A}}$ of the least favourable prior, $\pi_{!,A}$.
@@
On the flip side, however, this estimator $f_{\pi!,A}$ does not leverage the fact the low-MMSE nature of some prior, leading to suboptimal regret produced by $f_{\pi!,A}$. Indeed, we consider the priors generated by the neural prior on prior protocols, and the histogram of MMSEs as shown in Fig. 10(a). The MSE given by $f_{\pi!,[0,50]}$ on priors that are point masses as per Fig. 10(b) suggests that $f_{\pi!,[0,50]}$ is incapable of achieving low regrets on priors with low MMSEs.
@@ A.2 Training Priors
We now offer a more detailed description of the training priors.
Neural-generated: prior on priors. We sample the hidden mean parameter $\theta$ via the following: first, let $\mathcal{M}$ be classes of priors determined by some two-layer perceptron with a non-linear activation in-between. This is concretely defined as:
@@
$GELU,ReLU,SELU,CELU,SiLU,Tanh,TanhShrink.$
The parameter $\theta$ is then produced by sampling from a mixture of 4 priors in $\mathcal{M}$, and multiplied by $\theta_{\max}$ (or in the random $\theta_{\max}$ experiment as described in Section 5.1, each $\theta$ is then scaled differently).
Dirichlet process. Let the base distribution be defined as $H_{0}\triangleq\text{Unif}([0,h])$. Within each batch the elements $\theta_{1},\cdots,\theta_{s}$ are generated as follows:
$\theta_{j}=\begin{cases}\theta_{i}&\text{w.p. }\frac{j-1}{\alpha+j-1},\forall i=1,\cdots,j-1\\ x\sim H_{0}&\text{w.p. }\frac{\alpha}{\alpha+j-1}\end{cases}$
where $\alpha$ is a parameter that denotes how ‘close’ we are to iid generation ($\alpha=\infty$ essentially means we have iid). We use $\alpha=50$ for a sequence length of 512. Note that Dirichlet process implies that our data is not generated iid for each batch, so the Bayes estimator has to be estimated differently. We omit the calculation of this Bayes estimator.
@@ A.3 Why do we train using a mixture of two prior classes?
We consider the hypothesis: that our transformer trained under the mixture of the two priors is robust when evaluated under each of the priors. This can be verified via the following two tests: when evaluated on neural prior, is the performance (in terms of MSE) of the mixture-trained transformers closer to that of neural-trained ones as compared to the Dirichlet-trained ones? Similarly, when evaluated on Dirichlet prior, is the performance (in terms of MSE) of the mixture-trained transformers closer to that of Dirichlet-trained ones as compared to the neural-trained ones? Through the table of $T$-stat comparison done on the MSEs of 4096 seeds, we answer both these questions in the positive (the difference is especially obvious when evaluated on neural prior).
@@ B.1 Approximation of known empirical Bayes baselines
where $Q=YW_{Q},K=YW_{K},V=YW_{V}$, and $W_{Q},W_{K}\in\mathbb{R}^{(d+1)\times d_{k}}$. Let $Z=\text{Attention}(Q,K,V)$. We now design en encoding mechanism such that the representation after skip connection has the following:
@@
Define $D$ be a large number, $W_{Q}=I_{d+1}$ the $d$-dimensional identity matrix, $W_{V}=\begin{pmatrix}I_{d}&0\\ 0&0\\ \end{pmatrix}$ and $W_{K}\in\mathbb{R}^{d\times d}$ satisfying
@@
$(QK^{T})_{i,j}=\begin{cases}D&X_{i}=X_{j}=d\\ D+\sqrt{k}\log(X_{i}+1)&X_{j}=X_{i}+1\leq d\\ 0&\text{otherwise}\\ \end{cases}.$
@@
$M_{ij}=\begin{cases}\frac{1}{N(X_{i})+(X_{i}+1)N(X_{i}+1)}&X_{i}=X_{j}\leq d-1\\ \frac{X_{i}+1}{N(X_{i})+(X_{i}+1)N(X_{i}+1)}&X_{j}=X_{i}+1\leq d-1\\ 0&\text{otherwise}.\\ \end{cases}$
@@
Thus adding back $Y$ gives the desired output.
@@
Then $f$ is continuous, and $f(Z^{\prime})$ is indeed $\hat{\theta}_{\mathsf{Rob},d,M}$. Therefore by universal approximation theorem, there exists a multilayer perceptron that approximates $f$ within $[0,1]$, as desired. ∎
@@
Starting with the input tokens $(X_{1},\cdots,X_{n})$, we consider the token-wise embedding $Y_{i}=\mathsf{Sigmoid}(X_{i})$. Note that $\mathsf{Sigmoid}$ satisfies the assumption of $\varphi$ in Lemma B.1. Denote $p_{n}^{\mathsf{emp}}$ as the empirical distribution determined by $(X_{1},\cdots,X_{n})$. By Lemma B.1, the function $\hat{\theta}_{\mathsf{NPMLE},d}(p_{n}^{\mathsf{emp}},\cdot)$ can be continually extended (in the (weak, $\ell_{2}$) metric). Then [FdHP24], Theorem 1 says that there exists a transformers network $\Gamma$ that satisfies
@@
It now remains to consider the case where $x_{n}=x_{0}$ for all sufficiently large $n$; w.l.o.g. we may even assume $x_{n}=x_{0}$ for all $n$. If $x_{0}>d$ we are done since $\hat{\theta}_{\mathsf{NPMLE}}(p^{\mathsf{emp}},x_{0})=M$ for all $\pi$. Assume now that $x_{0}\leq d$. Recall that $\hat{\theta}_{\mathsf{NPMLE}}(p^{\mathsf{emp}},x_{0})=(x_{0}+1)\frac{f_{\hat{\pi}}(x_{0}+1)}{f_{\hat{\pi}}(x_{0})}$ where $\hat{\pi}$ is the prior estimated by NPMLE. Thus denoting $\hat{\pi}_{n}$ as NPMLE prior of $p^{\mathsf{emp}}_{n}$, for each $x$ it suffices to show that convergence of $f_{\hat{\pi}_{n}}(x)$. Now note that NPMLE also has the following equivalent form: $\hat{\pi}_{n}=\mathop{\rm argmin}_{Q}\text{KL}(p^{\mathsf{emp}}_{n}||f_{Q}$, and note that KL can be written in the following form (c.f. Assumption 1 of [JPW22]).
$\text{KL}(\pi_{1}||\pi_{2})=t(\pi_{1})+\sum_{x\geq 0}\ell(\pi_{1},\pi_{2})$ (3)
@@
$~{}\text{KL}(p_{0}||f_{\hat{\pi}_{n_{i}}})-\text{KL}(p_{0}||f_{Q_{0}})$
$\geq$ $~{}\text{KL}(p_{0}||f_{Q_{1}})-\text{KL}(p_{0}||f_{Q_{0}})\geq\delta$
@@
$~{}(\text{KL}(p_{n}^{\mathsf{emp}}||f_{Q})-\text{KL}(p_{n}^{\mathsf{emp}}||f_{Q^{\prime}}))$
-- $~{}(\text{KL}(p_{0}||f_{Q})-\text{KL}(p_{0}||f_{Q^{\prime}}))\to 0$
given that $p_{n}^{\text{emp}}\to p_{0}$ weakly and that $\text{KL}(p||f_{Q})-\text{KL}(p||f_{Q^{\prime}})=\sum_{y}p(y)\log\frac{f_{Q}^{\prime}}{f_{Q}}$, which is a contradiction. ∎
@@
Choose $d$ such that $\mathbb{P}[X>d]<\frac{\epsilon}{6\cdot\theta_{\max}^{2}}$. Note that there exists an $N$ such that for $n\geq N$, both the Robbins estimator [BGR13] and NPMLE [JPW22], Theorem 3 enjoy a minimax regret bounded by $\frac{\epsilon}{6}$ over the class $\mathcal{P}([0,\theta_{\max}])$. Now, by the previous two theorems, there exists a transformers model $\Gamma$ that can approximate either Robbins or NPMLE up to $\sqrt{\frac{\epsilon}{6}}$ precision uniformly for inputs up to $d$. Then we have
@@
$\leq 2(\frac{\epsilon}{6}+\mathbb{E}[(\hat{\theta}(X)-\Gamma(X))^{2}{\mathbf{1}_{\left\{{X\leq d}\right\}}}]$
$+\mathbb{E}[\theta_{\max}^{2}{\mathbf{1}_{\left\{{X>d}\right\}}}])$
$\leq 2(\frac{\epsilon}{6}+\frac{\epsilon}{6}+\frac{\epsilon}{6})$
@@
B.2 Identities on Worst Prior

### v2

@@ 7 Conclusion, limitation and future work
We have demonstrated the ability of transformers to learn EB-Poisson via in-context learning. This was done by evaluating pre-trained transformers on synthetic data of unseen distribution and sequence lengths, and compared against baselines like the NPMLE and other EB estimators. In this process, we showed that transformers can achieve decreasing regret as the sequence length increases. On the real datasets, we showed that these pre-trained transformers can outperform classical baselines in most cases.
Limitation. We humbly remark that we limited our focus to the 1-dimensional Poisson problem. Therefore, one future direction will be to extend our work to multi-dimensional input [JPTW23, Section 1.3], [JPW22, Section 6], and to other models like the normal-means model [JZ09]. We focused on the mean estimation problem, while a full scale posterior inference involves tasks like posterior sampling and uncertainty quantification. We believe that (probably larger) transformers could also achieve such tasks in the empirical Bayes setting. All these extensions, we believe, will expand the set of real-world applications, and can better demonstrate the power (or limitation) of transformers. While we established tools (Robbins and NPMLE approximation, and linear probe) to understand how transformers solve mean estimation in EB-Poisson setting, there are other aspects we have not discussed (e.g. training dynamics). These will be subject to future work.
@@ Acknowledgements
This work was supported in part by the MIT-IBM Watson AI Lab and the National Science Foundation under Grant No CCF-2131115. Anzo Teh was supported by a fellowship from the Eric and Wendy Schmidt Center at the Broad Institute. The authors acknowledge the MIT SuperCloud and Lincoln Laboratory Supercomputing Center for providing compute resources that have contributed to the research results reported within this paper.
@@ A.1 Worst-case prior and Gold-Standard Estimator
Let $A$ be a compact subset of $\mathbb{R}$. Then the worst-case prior $\pi_{!,A}$ is defined as
@@
One motivation for using the worst-case prior is that the Bayes estimator is considered the “gold standard” which minimizes the maximum possible MSE across all priors supported on $A$. A concrete statement can be found in the following lemma.
@@
Let $\hat{\theta}_{\pi}$ be the Bayes estimator of a given prior $\pi$, and let $A$ be any compact subset of the reals. Then the least favorable prior $\pi_{!,A}$ of $A$ satisfies the following:
@@
For any compact subset $A$ of the real numbers, we have
@@
achieved by the Bayes estimator $\hat{\theta}_{\pi_{!,A}}$ of the least favourable prior, $\pi_{!,A}$.
@@
Corollary A.3 says that among all the non-empirical Bayes estimators, $\hat{\theta}_{\pi_{!,A}}$ achieves the best minimax MSE loss among all priors supported on $A$. On the flip side, however, this estimator $\hat{\theta}_{\pi_{!,A}}$ does not leverage the fact the low-MMSE nature of some prior, leading to suboptimal regret produced by $\hat{\theta}_{\pi_{!,A}}$. In Fig. 5(a), we display the priors generated by the neural prior-on-priors and the histogram of MMSEs. The MSE given by $\hat{\theta}_{\pi_{!,[0,50]}}$ on priors that are point masses as per Fig. 5(b) suggests that $f_{\pi!,[0,50]}$ is incapable of achieving low regrets on priors with low MMSEs.
@@ A.2 Training priors
We now offer a more detailed description of the training priors. Recall in Section 3.3 we defined $\theta=\theta_{\mathsf{base}}\cdot\theta_{\max}$ with $\theta_{\mathsf{base}}\in[0,1]$. Here, we explain how $\theta_{\mathsf{base}}$ is generated according to both the neural-generated prior-on-priors, and the dirichlet process based on the uniform distribution. Recall that $\theta_{\max}$ is randomized at training phase according to how we described in Section 3.3, but nevertheless fixed during evaluation phase.
Neural-generated: prior-on-priors. We sample $\theta_{\mathsf{base}}\in[0,1]$ via the following: first, let $\mathcal{M}$ be classes of priors determined by some two-layer perceptron with a non-linear activation in-between. This is concretely defined as:
@@
$\{GELU,ReLU,SELU,CELU,SiLU,Tanh,TanhShrink\}.$
$\theta_{\mathsf{base}}$ is then produced by sampling from a mixture of 4 priors in $\mathcal{M}$.
Dirichlet process. Let the base distribution be defined as $H_{0}\triangleq\text{Unif}([0,1])$. Within each batch, elements $\theta_{\mathsf{base},1},\cdots,\theta_{\mathsf{base},n}$ are generated as follows:
$\theta_{\mathsf{base},j}=\begin{cases}\theta_{\mathsf{base},i}&\text{w.p. }\frac{j-1}{\alpha+j-1},\forall i=1,\cdots,j-1\\ \theta_{\mathsf{base}}\sim H_{0}&\text{w.p. }\frac{\alpha}{\alpha+j-1}\end{cases}$
where $\alpha$ is a parameter that denotes how ‘close’ we are to i.i.d. generation ($\alpha=\infty$ essentially means we have iid). We use $\alpha=50$ throughout (recall that the sequence length is 512). Note that Dirichlet process implies that our data is not generated i.i.d. for each batch, so the Bayes estimator has to be estimated differently. We omit the calculation of this Bayes estimator.
Mixture of two priors. We choose $\theta_{\mathsf{base}}$ from each of the two classes of prior with probability $\frac{1}{2}$, but each batch contains $\theta$’s drawn only from the same prior. This allows us to train our transformers to learn information about the prior $\pi$ that generates $\theta$ upon seeing sequence $(X_{1},X_{2},\cdots,X_{n})$.
@@ A.3 Why do we train using a mixture of two prior classes?
We consider the hypothesis: that our transformer trained under the mixture of the two priors is robust when evaluated under each of the priors. This can be verified via the following two tests: when evaluated on neural prior, is the performance (in terms of MSE) of the mixture-trained transformers closer to that of neural-trained ones as compared to the Dirichlet-trained ones? Similarly, when evaluated on Dirichlet prior, is the performance (in terms of MSE) of the mixture-trained transformers closer to that of Dirichlet-trained ones as compared to the neural-trained ones? By compring the MSEs of 4096 seeds at evaluation stage at Table 2, we answer both these questions in the positive (the difference is especially obvious when evaluated on neural prior).
@@ B.1 Approximation of known empirical Bayes baselines
where $Q=YW_{Q},K=YW_{K},V=YW_{V}$, and $W_{Q},W_{K}\in\mathbb{R}^{(d+1)\times(d+1)}$. Let $Z=\text{Attention}(Q,K,V)$. We now design an encoding mechanism such that the representation after skip connection has the following:
@@
Define $D$ be a large number, $W_{Q}=I_{d+1}$ the $d$-dimensional identity matrix, $W_{V}=\begin{pmatrix}I_{d}&0\\ 0&0\\ \end{pmatrix}$ and $W_{K}\in\mathbb{R}^{(d+1)\times(d+1)}$ satisfying
@@
$(QK^{T})_{i,j}=\begin{cases}D&X_{i}=X_{j}\leq d\\ D+\sqrt{k}\log(X_{i}+1)&X_{j}=X_{i}+1\leq d\\ 0&\text{otherwise}\\ \end{cases}.$
@@
$M_{ij}=\begin{cases}\frac{1}{N(X_{i})+(X_{i}+1)N(X_{i}+1)}&X_{i}=X_{j}\leq d-1\\ \frac{X_{i}+1}{N(X_{i})+(X_{i}+1)N(X_{i}+1)}&X_{j}=X_{i}+1\leq d\\ 0&\text{otherwise}.\\ \end{cases}$
@@
Thus, adding back $Y$ gives the desired output.
@@
Then $f$ is continuous, and $f(Z^{\prime})$ is indeed $\hat{\theta}_{\mathsf{Rob},d,M}$. In addition, this $f$ is $(1+M)^{2}$-Lipchitz, and therefore by the results in [Mha96, Theorem 2.1], there exists a two-layered network with $O((1+M)^{2}\epsilon^{-1})$ hidden neurons that approximates $f$ within $[0,1]$ with error at most $\epsilon$, as desired. ∎
Remark B.1.
We note that Robbins approximation using linear attention is also possible, using the attention defined in [KVPF20, (4)]. Specifically, We embed our inputs representation $X\in\mathbb{R}^{n}$ into one-hot vector $Y\in\mathbb{R}^{n\times(d+1)}$ where $Y_{i}=e_{X_{i}+1}$ if $X_{i}<d$ and $e_{d+1}$ if $X_{i}\geq d$ (instead of 0 to prevent the denominator in [KVPF20, (4)] for normalization to be 0). We choose $\phi$ the identity function, $W_{Q}=I_{d+1}$ the $d+1$-dimensional identity matrix, $W_{V}=\begin{pmatrix}I_{d}&0\\ 0&0\\ \end{pmatrix}$, and $W_{K}$ the following tridiagonal matrix:
$(W_{K})_{ij}=\begin{cases}1&i=j\\ j&j=1\leq i-1\leq d-1\\ 0&\text{otherwise}\\ \end{cases}$
This results in the following:
$Q_{i}=\begin{cases}e_{X_{i}+1}&X_{i}<d\\ e_{d+1}&\text{otherwise}\\ \end{cases}\qquad K_{j}=\begin{cases}e_{X_{j}+1}+X_{j}e_{X_{j}}&X_{j}<d\\ e_{d+1}&\text{otherwise}\\ \end{cases}\qquad V_{j}=\begin{cases}e_{X_{j}+1}&X_{j}<d\\ 0&\text{otherwise}\\ \end{cases}$
and will therefore yield the same attention output after normalization.
@@
Starting with the input tokens $(X_{1},\cdots,X_{n})$, we consider the token-wise embedding $Y_{i}=\mathsf{Sigmoid}(X_{i})$. Note that $\mathsf{Sigmoid}$ satisfies the assumption of $\varphi$ in Lemma B.2. Denote $p_{n}^{\mathsf{emp}}$ as the empirical distribution determined by $(X_{1},\cdots,X_{n})$. By Lemma B.2, the function $\hat{\theta}_{\mathsf{NPMLE},d}(p_{n}^{\mathsf{emp}},\cdot)$ can be continuously extended (in the (weak, $\ell_{2}$) metric). Then [FdHP24], Theorem 1 says that there exists a transformers network $\Gamma$ that satisfies
@@
It now remains to consider the case where $x_{n}=x_{0}$ for all sufficiently large $n$; w.l.o.g. we may even assume $x_{n}=x_{0}$ for all $n$. If $x_{0}>d$ we are done since $\hat{\theta}_{\mathsf{NPMLE}}(p^{\mathsf{emp}},x_{0})=M$ for all $\pi$. Assume now that $x_{0}\leq d$. Recall that $\hat{\theta}_{\mathsf{NPMLE}}(p^{\mathsf{emp}},x_{0})=(x_{0}+1)\frac{f_{\hat{\pi}}(x_{0}+1)}{f_{\hat{\pi}}(x_{0})}$ where $\hat{\pi}$ is the prior estimated by NPMLE. Thus denoting $\hat{\pi}_{n}$ as NPMLE prior of $p^{\mathsf{emp}}_{n}$, for each $x$ it suffices to show that convergence of $f_{\hat{\pi}_{n}}(x)$. Now note that NPMLE also has the following equivalent form: $\hat{\pi}_{n}=\mathop{\rm argmin}_{Q}D(p^{\mathsf{emp}}_{n}||f_{Q})$, where $D$ denotes the KL divergence. Note that $D$ can be written in the following form (c.f. Assumption 1 of [JPW22]).
$D(\pi_{1}||\pi_{2})=t(\pi_{1})+\sum_{x\geq 0}\ell(\pi_{1},\pi_{2})$ (4)
@@
$~{}D(p_{0}||f_{\hat{\pi}_{n_{i}}})-D(p_{0}||f_{Q_{0}})\geq D(p_{0}||f_{Q_{1}})-D(p_{0}||f_{Q_{0}})\geq\delta$
@@
$(D(p_{n}^{\mathsf{emp}}||f_{Q})-D(p_{n}^{\mathsf{emp}}||f_{Q^{\prime}}))-(D(p_{0}||f_{Q})-D(p_{0}||f_{Q^{\prime}}))\to 0$
given that $p_{n}^{\text{emp}}\to p_{0}$ weakly and that $D(p||f_{Q})-D(p||f_{Q^{\prime}})=\sum_{y}p(y)\log\frac{f_{Q}^{\prime}}{f_{Q}}$, which is a contradiction. ∎
@@
Choose $d$ such that $\mathbb{P}[X>d]<\frac{\epsilon}{6\cdot\theta_{\max}^{2}}$. By the properties of the Poisson distribution (e.g. [PW25, (15.19)] on Poisson-Binomial distribution), for each $x>\theta_{\max}$ we have
$\mathbb{P}[X\geq x]\leq\sup_{0\leq\theta\leq\theta_{\max}}\frac{(e\theta)^{x}e^{-\theta}}{x^{x}}\leq\frac{(e\theta_{\max})^{x}e^{-\theta_{\max}}}{x^{x}}$
Thus there exists some $d\in O(\log(\frac{1}{\epsilon})+\log\theta_{\max})$ that fulfills this condition.
Next, with sequence length $n$, both the Robbins estimator [PW21, Theorem 2] and NPMLE [JPW22, Theorem 3],enjoy a minimax regret of $O_{\theta_{\max}}(\frac{1}{n}(\frac{\log n}{\log\log n})^{2})$. This means we can choose $N=O(\frac{1}{\epsilon}\log^{2}\frac{1}{\epsilon})$ such that for all $n\geq N$, the minimax regret over the class $\mathcal{P}([0,\theta_{\max}])$ is bounded by $\frac{\epsilon}{6}$. Now, by the previous two theorems, for each $M$, there exists a transformer model $\Gamma$ that can approximate either Robbins or NPMLE (clipped at $M$) up to $\sqrt{\frac{\epsilon}{6}}$ precision uniformly for inputs up to $d$. Given that $\theta\leq\theta_{\max}$, we can choose the threshold $M=\theta_{\max}$ in both cases without increasing the regret. Then we have
@@
$\leq 2\left(\frac{\epsilon}{6}+\mathbb{E}[(\hat{\theta}(X)-\Gamma(X))^{2}{\mathbf{1}_{\left\{{X\leq d}\right\}}}]+\mathbb{E}[\theta_{\max}^{2}{\mathbf{1}_{\left\{{X>d}\right\}}}]\right)$
$\leq 2\left(\frac{\epsilon}{6}+\frac{\epsilon}{6}+\frac{\epsilon}{6}\right)$
@@
Finally, to bound the model complexity of a transformer needed for such approximation, we follow the Robbins approximation as per Theorem 4.1. Here, we used an encoder with embedding dimension of $O(\log(\epsilon^{-1}))$, and decoder an multi-layer perceptrons with $O((1+\theta_{\max})^{2}\epsilon^{-1})$ layers and width 2. This gives a total of $O((1+\theta_{\max})^{2}\epsilon^{-1}+\log^{2}(\epsilon^{-1}))=O_{\theta_{\max}}(\epsilon^{-1})$ parameters. ∎
B.2 Identities on worst-case prior